# Preference Optimization with Multi-Sample Comparisons

### v1

@@ 1 Introduction
Generative models, particularly large language models (LLMs) (Achiam et al., 2023; Meta AI, 2024; Bai et al., 2023; Bi et al., 2024; Gemini et al., 2023; Anthropic, 2024) and diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Rombach et al., 2022; Podell et al., 2023), hold the tremendous promise to transform numerous industries by automating complex tasks, enhancing creativity, and personalizing user experiences at an unprecedented scale (Eloundou et al., 2023). These models achieve their capabilities through extensive pretraining on large-scale datasets, followed by post-training to unlock the capabilities (i.e., the superficial alignment hypothesis) (Wei et al., 2022; Tay et al., 2022; Zhou et al., 2024a). For post-training stage, reinforcement learning from human feedbacks (RLHF) and direct alignment from preference methods (DAP) have been crucial for the success of both LLMs (Ouyang et al., 2022; Rafailov et al., 2024; Azar et al., 2024) and diffusion models (Black et al., 2023; Wallace et al., 2023; Yang et al., 2024).
@@
In this work, we introduce Multi-sample Direct Preference Optimization (mDPO) and Multi-sample Identity Preference Optimization (mIPO), which are extensions of the prior DAP methods DPO (Rafailov et al., 2024) and IPO (Azar et al., 2024) [^1] Our framework can also be extended to other preference optimization algorithms.. Unlike their predecessors, which rely on single-sample comparisons, mDPO and mIPO utilize multi-sample comparisons to better capture group-wise or distributional characteristics. Our experiments involve fine-tuning language models to generate random numbers in a calibrated manner and enhancing the diversity of genres in creative story writing. Additionally, for diffusion models, we demonstrate a significant reduction in gender and race biases compared to the traditional single-sample comparison approach. Furthermore, we present evidence that multi-sample comparison offers a more robust optimization method in the presence of label noise in preference data, such as synthetic data that lacks human labeling but possesses knowledge of the overall quality between two models.
@@ 2 Related Works
While finalizing this manuscript for submission to arXiv, we became aware of a concurrent study by Li et al. (2024). Both works independently address the formulation of multi-sample comparison. Our contributions differ in several key aspects, such as the introduction of additional algorithms, distinct experimental foci, and differing applications. Notably, our primary focus is on language models, with an emphasis on providing a low-variance and potentially unbiased estimator. In contrast, their work centers around text-to-image tasks and does not include the IPO-based methods we derived.
@@ 3 Preliminaries
Supervised Finetuning. After the language model has been pretrained on extensive datasets, the next step is typically supervised finetuning (SFT). This process aims to refine the model’s predictions, preparing it for subsequent alignment stages. During SFT, the pretrained language model is finetuned using a supervised dataset $\mathcal{D}=\{(x,y)_{i}\}_{i=1}^{N}$, where $x$ represents the input and $y$ denotes the target. The objective of SFT is to minimize the negative log-likelihood,
@@
Direct Alignment from Preference. For single-sample comparison setting, the dataset consists of $\{(x,y_{w},y_{l})_{i}\}_{i=1}^{N}$, where $x$ denotes the prompt or the input to the model, $y_{w}$ is the preferred response over $y_{l}$. Under the Bradley-Terry model, the likelihood of $y_{w}$ is preferred to $y_{l}$ is computed by
@@
where $r(y,x)$ is the scalar reward of answering $y$ when given $x$. Direct preference optimization (DPO) (Rafailov et al., 2024) establishes an analytical form between the reward function and the policy or the language model, i.e., the language model represents an implicit reward. Specifically, under DPO, the implicit reward can be computed by
@@
Using the implicit reward, DPO optimizes the policy $\pi_{\bm{\theta}}$ by minimizing the negative log-likelihood on the offline preference dataset,
@@
Although the Bradley-Terry model is based on the assumption of Gumbel noise in the reward function, a choice that is well-suited for discrete random variables, it has certain limitations, such as overfitting, as noted by Azar et al. (2024). Specifically, Identity Preference Optimization (IPO) (Azar et al., 2024) addresses these limitations by bypassing the Bradley-Terry model for preference modeling. IPO aims to mitigate the issue of “overfitting” in preference datasets by replacing the sigmoid function with the squared distance,
$\mathcal{L}_{\mathrm{IPO}}(\pi_{\bm{\theta}},\mathcal{D})=\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\frac{\pi_{\bm{\theta}}(y_{w}|x)}{\pi_{\textnormal{ref}}(y_{l}|x)}-\log\frac{\pi_{\bm{\theta}}(y_{l}|x)}{\pi_{\textnormal{ref}}(y_{l}|x)}-\frac{\tau^{-1}}{2}\right]^{2},$
@@ 4.1 Multi-sample Preference Optimization
Thus, we generalize the preference model from singleton comparison to multi-sample comparison, where the preference is assigned at the multi-sample level instead of on a singleton-level. Let $\mathcal{G}_{w}$ and $\mathcal{G}_{l}$ denote the multiple samples sampled from the model’s conditional distribution [^2] The definition of a group $\mathcal{G}$ can be generalized to distributions. Thus, $\mathcal{G}$ is equivalent to the predictive distribution of the language model. when given the prompt $x$, then the likelihood of $\mathcal{G}_{w}$ is preferred over $\mathcal{G}_{l}$ is defined as
@@
where the function $\Phi$ follows the definition in the IPO paper (Azar et al., 2024), which can be e.g., the sigmoid function (recovers the Bradley-Terry model). Given the reward function $r(\cdot,\cdot)$, the goal of RLHF is to maximize the reward under reverse KL constraints. This can be captured by the following optimization objective,
@@
Generalizing the DPO’s derivation, we obtain the following equation for the reward function,
@@
$\mathcal{L}_{\mathrm{mDPO}}$ $=\mathbb{E}_{(x,\mathcal{G}_{w},\mathcal{G}_{l})\sim\mathcal{D}}\left[-\log\sigma\left(\beta{\color[rgb]{0.84375,0.1875,0.15234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.84375,0.1875,0.15234375}\mathbb{E}_{y_{w}\sim\mathcal{G}_{w}}}\left[\log\frac{\pi_{\bm{\theta}}(y_{w}|x)}{\pi_{\textnormal{ref}}(y_{w}|x)}\right]-\beta{\color[rgb]{0.84375,0.1875,0.15234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.84375,0.1875,0.15234375}\mathbb{E}_{y_{l}\sim\mathcal{G}_{l}}}\left[\log\frac{\pi_{\bm{\theta}}(y_{l}|x)}{\pi_{\textnormal{ref}}(y_{l}|x)}\right]\right)\right].$
@@
$\mathcal{L}_{\mathrm{mIPO}}$ $=\mathbb{E}_{(x,\mathcal{G}_{w},\mathcal{G}_{l})\sim\mathcal{D}}\left[\left({\color[rgb]{0.84375,0.1875,0.15234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.84375,0.1875,0.15234375}\mathbb{E}_{y_{w}\sim\mathcal{G}_{w}}}\left[\log\frac{\pi_{\bm{\theta}}(y_{w}|x)}{\pi_{\textnormal{ref}}(y_{w}|x)}\right]-{\color[rgb]{0.84375,0.1875,0.15234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.84375,0.1875,0.15234375}\mathbb{E}_{y_{l}\sim\mathcal{G}_{l}}}\left[\log\frac{\pi_{\bm{\theta}}(y_{l}|x)}{\pi_{\textnormal{ref}}(y_{l}|x)}\right]-\frac{\tau^{-1}}{2}\right)^{2}\right].$
The main difference (highlighted above in red) from their original objectives is that the implicit reward for a single sample is replaced with the averaged implicit reward for a group of samples. We use the notation $y\sim\mathcal{G}$ to account for the general case where $\mathcal{G}$ represents a distribution rather than merely a finite collection of samples.
@@ 4.2 Stochastic Estimator for Efficient Optimization
$\mathbb{E}_{(x,\{y_{w,i}\}_{i=1}^{k},\{y_{l,i}\}_{i=1}^{k})\sim\mathcal{D}}\left[\left(\frac{1}{k}\sum_{i=1}^{k}\log\frac{\pi_{\bm{\theta}}(y_{w,i}|x)}{\pi_{\textnormal{ref}}(y_{w,i}|x)}-\frac{1}{k}\sum_{j=1}^{k}\log\frac{\pi_{\bm{\theta}}(y_{l,j}|x)}{\pi_{\textnormal{ref}}(y_{l,j}|x)}-\frac{\tau^{-1}}{2}\right)^{2}-\frac{\hat{\sigma}_{w}^{2}}{k}-\frac{\hat{\sigma}_{l}^{2}}{k}\right].$
@@
$\mathbb{E}_{(x,\{y_{w,i}\}_{i=1}^{k},\{y_{l,i}\}_{i=1}^{k})\sim\mathcal{D}}\left[-\log\sigma\left(\frac{\beta}{k}\sum_{i=1}^{k}\log\frac{\pi_{\bm{\theta}}(y_{w,i}|x)}{\pi_{\textnormal{ref}}(y_{w,i}|x)}-\frac{\beta}{k}\sum_{j=1}^{k}\log\frac{\pi_{\bm{\theta}}(y_{l,j}|x)}{\pi_{\textnormal{ref}}(y_{l,j}|x)}\right)\right].$
@@ 5.1 Random Number Generator (RNG)
Figure 3: Dist. comparison.
@@
Preference data construction and finetuning. Our initial experiments focus on generating integers uniformly at random within an interval $[a,b]$, where $a$ and $b$ are integers. To create the dataset, we sample $a$ randomly from $[0,1000]$ and the interval gap $m$ from $[5,10]$, with $b$ defined as $a+m$. The preferred response group is based on a uniform distribution of numbers, while non-uniform distributions are classified as rejected. We generated approximately 3,000 paired groups for training. For testing, we sample $a$ and $b$ from $[0,1000]$ with $a<b$ but do not impose the same gap constraints as in the training set to evaluate generalizability. We implemented the multi-sample versions of IPO and DPO, named mIPO and mDPO, respectively. To stabilize training, we add a negative log-likelihood [^3] Without this, the language model tends to generate numbers beyond the specified range. We discuss this further from a constrained optimization perspective. to the objective of both the multi-sample and original versions of IPO and DPO, shown to be effective in reasoning tasks (Pang et al., 2024). For finetuning, we use the Llama 3-8B instruct model with LoRA, rank 64, and $\alpha=128$. More training details are in Appendix A.1.
@@ 5.2 Controlled Debiasing for Image Generation
Diffusion models (Sohl-Dickstein et al., 2015) have emerged as a powerful workhorse for image generation (Ho et al., 2020; Rombach et al., 2022; Dai et al., 2023; Esser et al., 2024). Despite their success, the generated images often reflect strong biases and stereotypes due to imbalances in the training data (Ananya, 2024). These biases can result in harmful societal stereotypes and limit the diversity of generated content. To illustrate this point, we visualize the distribution of generated images for different occupations and highlight the discrimination in race and biases in Figure 11 in the Appendix using Stable diffusion 1.5 (Rombach et al., 2022).
To address the generation biases, we extend the diffusion DPO objective proposed in Wallace et al. (2023) from single-sample comparison to multi-sample comparison in a similar way as we discussed in Section 4. For diffusion models, the mDPO objective can be formulated as [^4] It’s more appropriate to interpret $r(\mathbf{\epsilon},{\bm{x}};\theta)$ as a cost rather than a reward. Consequently, the formulation uses $-\beta$ instead of $\beta$, as done in the original DPO.
$\mathcal{L}_{\text{mDPO}}^{\text{diff}}(\epsilon_{\bm{\theta}},\mathcal{D})=\mathbb{E}\left[-\log\sigma\left(-\beta T\omega(\lambda_{t})\cdot\left(\mathbb{E}_{\mathcal{G}_{w}}\left[r(\mathbf{\epsilon}^{w},{\bm{x}}_{t}^{w},t;{\bm{\theta}})\right]-\mathbb{E}_{\mathcal{G}_{l}}\left[r(\mathbf{\epsilon}^{l},{\bm{x}}_{t}^{l},t;{\bm{\theta}})\right]\right)\right)\right],$
@@
Dataset construction and finetuning. To construct the dataset, we first collect $50$ occupations, and take $80\%$ of the occupations for generating training data and the remaining $20\%$ for evaluation. For each occupation, we generate $6$ images with varied races in {black, white, aisan} and genders in {male, female}, which form our chosen group. For the rejected group, we use the images generated from the original Stable diffusion 1.5 via API [^5] https://stablediffusionapi.com/docs/category/stable-diffusion-api. Our training code is adapted from the original implementation by Wallace et al. (2023). More details on the training can be found in Appendix A.2.
Evaluation metric and results. To measure the generation quality, we used the Simpson Diversity Index (Simpson, 1949), which focuses more on the dominance and even distribution of species. To compute it, we use $D=1-\sum_{i}(n_{i}/N)^{2}$, where $n_{i}$ is the number of instance falls in $i_{\text{th}}$ category (e.g., for gender, the categories we consider are {male, female}), and $N=\sum_{i}n_{i}$. In general, the higher the value, the better the diversity. We compared three methods, SFT, DPO (equivalent to $k=1$ for mDPO), and mDPO ($k=6$). The results are presented in Table 2.
@@
We observe that mDPO performs the best in both metrics, which improves the Simpson Diversity Index signifacantly over the original SD 1.5, but DPO did not improve over SFT and even perform worse than SFT on Gender. For qualitative comparison, we visualized $10$ images generated by the original diffusion model and the diffusion model after mDPO finetuning in Figure 8 using the testing prompt. We observe that the finetuned model generates more balanced and diversified images in terms of race and gender than the baseline.

### v2

@@ 1 Introduction
Generative models, particularly large language models (LLMs) (Achiam et al., 2023; Meta AI, 2024; Bai et al., 2023; Bi et al., 2024; Gemini et al., 2023; Anthropic, 2024) and diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Rombach et al., 2022; Podell et al., 2023), hold the tremendous promise to transform numerous industries by automating complex tasks, enhancing creativity, and personalizing user experiences at an unprecedented scale (Eloundou et al., 2023). These models achieve their capabilities through extensive pretraining on large-scale datasets, followed by post-training to unlock the capabilities (i.e., the superficial alignment hypothesis) (Wei et al., 2022; Tay et al., 2022; Zhou et al., 2024a). For post-training stage, reinforcement learning from human feedback (RLHF) and direct alignment from preference methods (DAP) have been crucial for the success of both LLMs (Rafailov et al., 2024; Azar et al., 2024) and diffusion models (Black et al., 2023; Wallace et al., 2023; Yang et al., 2024).
@@
In this work, we introduce Multi-sample Direct Preference Optimization (mDPO) and Multi-sample Identity Preference Optimization (mIPO), which are extensions of the prior DAP methods–DPO (Rafailov et al., 2024) and IPO (Azar et al., 2024) [^1] Our framework can also be extended to other preference optimization algorithms.. Unlike their predecessors, which rely on single-sample comparisons, mDPO and mIPO utilize multi-sample comparisons to better capture group-wise or distributional characteristics. Our experiments involve fine-tuning language models to generate random numbers in a calibrated manner and enhancing the diversity of genres in creative story writing. Additionally, for diffusion models, we demonstrate a significant reduction in gender and race biases compared to the traditional single-sample comparison approach. Furthermore, we present remarkable results where mDPO and mIPO can iteratively improve the model quality by only contrasting multiple responses generated by a stronger and weaker model without using precise pairwise comparisons.
@@ 2 Related Works
While finalizing this manuscript for submission to arXiv, we became aware of a related study by Li et al. (2024). Both works independently address the formulation of multi-sample comparison. Our contributions differ in several key aspects, such as the introduction of additional algorithms, distinct experimental foci, and differing applications. Notably, our primary focus is on language models, with an emphasis on providing a low-variance and potentially unbiased estimator. In contrast, their work centers around text-to-image tasks and does not include the IPO-based methods we derived.
@@ 3 Preliminaries
Supervised Finetuning. After the language model has been pretrained extensively on large corpora, the next step is typically supervised finetuning (SFT). This process aims to teach the model the dialog formats and refine its predictions to interact more naturally with human and prepare it for the subsequent alignment stages. Specifically, the pretrained language model is finetuned using a supervised dataset $\mathcal{D}=\{(x,y)_{i}\}_{i=1}^{N}$, where $x$ represents the input and $y$ denotes the target. The objective of SFT is to minimize the negative log-likelihood,
@@
Direct Alignment from Preference.
For single-sample comparison setting, the dataset consists of $\{(x,y_{w},y_{l})_{i}\}_{i=1}^{N}$, where $x$ denotes the prompt or the input to the model, $y_{w}$ is the preferred response over $y_{l}$. Under the Bradley-Terry model, the likelihood of $y_{w}$ is preferred to $y_{l}$ is computed as
@@
where $r(y,x)$ is the scalar reward of answering $y$ given $x$. Direct preference optimization (DPO) (Rafailov et al., 2024) establishes an analytical form between the reward function and the policy language model, where the language model’s prediction represents an implicit reward. Specifically, under DPO, the implicit reward can be computed by
@@
Using this implicit reward, DPO optimizes the policy $\pi_{\bm{\theta}}$ by minimizing the negative log-likelihood on the offline preference dataset,
@@
The Bradley-Terry model is based on the assumption of Gumbel noise in the reward function, which is well-suited for discrete random variables. However, it has certain limitations such as overfitting, as noted by Azar et al. (2024). Specifically, Identity Preference Optimization (IPO) (Azar et al., 2024) addresses these limitations by bypassing the Bradley-Terry model for preference modeling. IPO aims to mitigate the issue of “overfitting" in preference datasets by replacing the sigmoid function with the squared distance,
$\mathcal{L}_{\mathrm{IPO}}(\pi_{\bm{\theta}},\mathcal{D})=\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\bigg{[}\log\frac{\pi_{\bm{\theta}}(y_{w}|x)}{\pi_{\textnormal{ref}}(y_{l}|x)}-\log\frac{\pi_{\bm{\theta}}(y_{l}|x)}{\pi_{\textnormal{ref}}(y_{l}|x)}-\frac{\tau^{-1}}{2}\bigg{]}^{2},$
@@ 4.1 Multi-sample Preference Optimization
We generalize the preference model from singleton comparison to multi-sample comparison, where the preference is assigned at the multi-sample level instead of on a singleton-level. Let $\mathcal{G}_{w}$ and $\mathcal{G}_{l}$ denote two group of samples drawn from some models’ conditional distributions when the prompt $x$ is given. [^2] The definition of a group $\mathcal{G}$ can be generalized to distributions. Thus, $\mathcal{G}$ is equivalent to the predictive distribution of the language model. Following the definitions from Azar et al. (2024), we define the likelihood of $\mathcal{G}_{w}$ preferred over $\mathcal{G}_{l}$ as
@@
where $\Phi$ can be a sigmoid function (recovering the Bradley-Terry model). Given the reward function $r(\cdot,\cdot)$, the goal of RLHF is to maximize the reward under reverse KL constraints. This can be captured by the following optimization objective,
@@
Generalizing the derivation of DPO, we obtain the following equation for the reward function,
@@
$\mathcal{L}_{\mathrm{mDPO}}=\mathbb{E}_{(x,\mathcal{G}_{w},\mathcal{G}_{l})\sim\mathcal{D}}\bigg{[}-\log\sigma\bigg{(}\beta{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbb{E}_{y_{w}\sim\mathcal{G}_{w}}}\left[\log\frac{\pi_{\bm{\theta}}(y_{w}|x)}{\pi_{\textnormal{ref}}(y_{w}|x)}\right]-\beta{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbb{E}_{y_{l}\sim\mathcal{G}_{l}}}\left[\log\frac{\pi_{\bm{\theta}}(y_{l}|x)}{\pi_{\textnormal{ref}}(y_{l}|x)}\right]\bigg{)}\bigg{]}.$ (1)
@@
$\mathcal{L}_{\mathrm{mIPO}}=\mathbb{E}_{(x,\mathcal{G}_{w},\mathcal{G}_{l})\sim\mathcal{D}}\bigg{[}\bigg{(}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbb{E}_{y_{w}\sim\mathcal{G}_{w}}}\left[\log\frac{\pi_{\bm{\theta}}(y_{w}|x)}{\pi_{\textnormal{ref}}(y_{w}|x)}\right]-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbb{E}_{y_{l}\sim\mathcal{G}_{l}}}\left[\log\frac{\pi_{\bm{\theta}}(y_{l}|x)}{\pi_{\textnormal{ref}}(y_{l}|x)}\right]-\frac{\tau^{-1}}{2}\bigg{)}^{2}\bigg{]}.$ (2)
The main difference (highlighted in Eq. (1) and Eq. (2) with red) from their original objectives is that the implicit reward for a single sample is replaced with the expectation of the implicit reward from the distribution $\mathcal{G}_{w}$ or $\mathcal{G}_{l}$. In practice, we use a finite set of examples collected from $\mathcal{G}_{w}$ and $\mathcal{G}_{l}$ and take the average of their implicit rewards to approximate the expectations. In the next section, we analyze the bias and variance of this approximation.
@@ 4.2 Stochastic Estimator for Efficient Optimization
$\mathbb{E}_{(x,\{y_{w,i}\}_{i=1}^{k},\{y_{l,i}\}_{i=1}^{k})\sim\mathcal{D}}\Bigg{[}\Bigg{(}\frac{1}{k}\sum_{i=1}^{k}\log\frac{\pi_{\bm{\theta}}(y_{w,i}|x)}{\pi_{\textnormal{ref}}(y_{w,i}|x)}-\frac{1}{k}\sum_{j=1}^{k}\log\frac{\pi_{\bm{\theta}}(y_{l,j}|x)}{\pi_{\textnormal{ref}}(y_{l,j}|x)}-\frac{\tau^{-1}}{2}\Bigg{)}^{2}-\frac{\hat{\sigma}_{w}^{2}}{k}-\frac{\hat{\sigma}_{l}^{2}}{k}\Bigg{]}.$
@@
$\mathbb{E}_{(x,\{y_{w,i}\}_{i=1}^{k},\{y_{l,i}\}_{i=1}^{k})\sim\mathcal{D}}\Bigg{[}-\log\sigma\Bigg{(}\frac{\beta}{k}\sum_{i=1}^{k}\log\frac{\pi_{\bm{\theta}}(y_{w,i}|x)}{\pi_{\textnormal{ref}}(y_{w,i}|x)}-\frac{\beta}{k}\sum_{j=1}^{k}\log\frac{\pi_{\bm{\theta}}(y_{l,j}|x)}{\pi_{\textnormal{ref}}(y_{l,j}|x)}\Bigg{)}\Bigg{]}.$
@@ 5.1 Random Number Generator (RNG)
Figure 3: Distance comparison.
@@
Preference data construction and finetuning. Our initial experiments focus on generating integers uniformly at random within an interval $[a,b]$, where $a$ and $b$ are integers. To create the dataset, we sample $a$ randomly from $[0,1000]$ and the interval gap $m$ from $[5,10]$, with $b$ defined as $a+m$. The preferred response group is based on a uniform distribution of numbers, while non-uniform distributions are classified as rejected. We generated approximately 3,000 paired groups for training. For testing, we sample $a$ and $b$ from $[0,1000]$ with $a<b$ but do not impose the same gap constraints as in the training set to evaluate generalizability. We implemented the multi-sample versions of IPO and DPO, named mIPO and mDPO, respectively. To stabilize training, we add a negative log-likelihood [^3] Without this, the language model tends to generate numbers beyond the specified range. We discuss this further from a constrained optimization perspective. to the objective of both the multi-sample and original versions of IPO and DPO, shown to be effective in reasoning tasks (Pang et al., 2024). For finetuning, we use the Llama 3-8B instruct model with LoRA (r=64, $\alpha=128$). More details are in Appendix 7.1.
@@ 5.2 Controlled Debiasing for Image Generation
Diffusion models (Sohl-Dickstein et al., 2015) have emerged as a powerful workhorse for image generation (Ho et al., 2020; Rombach et al., 2022; Dai et al., 2023; Esser et al., 2024). Despite their success, the generated images often reflect strong biases and stereotypes due to imbalances in the training data (Ananya, 2024). These biases can result in harmful societal stereotypes and limit the diversity of generated content. To illustrate this point, we visualize the distribution of generated images for different occupations and highlight the discrimination in race and biases in Figure 10 in the Appendix using Stable diffusion 1.5 (Rombach et al., 2022). To address the generation biases, we extend the diffusion DPO objective proposed in Wallace et al. (2023) from single-sample comparison to multi-sample comparison in a similar way as we discussed in Section 4. For diffusion models, the mDPO objective can be formulated as [^4] It’s more appropriate to interpret $r(\mathbf{\epsilon},{\bm{x}};\theta)$ as a cost rather than a reward. Consequently, the formulation uses $-\beta$ instead of $\beta$, as done in the original DPO.
$\mathcal{L}_{\text{mDPO}}^{\text{diff}}(\epsilon_{\bm{\theta}},\mathcal{D})=\mathbb{E}\big{[}-\log\sigma\big{(}-\beta T\omega(\lambda_{t})\cdot\big{(}\mathbb{E}_{\mathcal{G}_{w}}\left[r(\mathbf{\epsilon}^{w},{\bm{x}}_{t}^{w},t;{\bm{\theta}})\right]-\mathbb{E}_{\mathcal{G}_{l}}\left[r(\mathbf{\epsilon}^{l},{\bm{x}}_{t}^{l},t;{\bm{\theta}})\right]\big{)}\big{)}\big{]},$ (3)
@@
Dataset construction and finetuning. To construct the dataset, we first collect $50$ occupations, and take $80\%$ of the occupations for generating training data and the remaining $20\%$ for evaluation. For each occupation, we generate $6$ images with varied races in {black, white, aisan} and genders in {male, female}, which form our chosen group. For the rejected group, we use the images generated from the original Stable diffusion 1.5 via API [^5] https://stablediffusionapi.com/docs/category/stable-diffusion-api. Our training code is adapted from the original implementation by Wallace et al. (2023). More details on the training can be found in Appendix 7.2.
Evaluation metric and results. To measure the generation quality, we used the Simpson Diversity Index (Simpson, 1949), which focuses more on the dominance and even distribution of species. To compute it, we use $D=1-\sum_{i}(n_{i}/N)^{2}$, where $n_{i}$ is the number of instance falls in $i_{\text{th}}$ category (e.g., for gender, the categories we consider are {male, female}), and $N=\sum_{i}n_{i}$.
@@
In general, the higher the value, the better the diversity. We compared three methods, SFT, DPO (equivalent to $k=1$ for mDPO), and mDPO ($k=6$). The results are presented in Table 2. We observe that mDPO performs the best in both metrics, improving the Simpson Diversity Index significantly over the original SD 1.5. For qualitative comparison, we visualized $10$ images generated by the original diffusion model and the diffusion model after mDPO finetuning in Appendix 8.1 using the testing prompt. We observe that the finetuned model generates more balanced and diversified images in terms of race and gender than the baseline.